OpenStudy (anonymous):

If G(x) = (x + 3)2, and F(x) = x2 + 7x, what is the function G(F(x))? A. (x2 + 7x + 3)2 B. (x + 3)4 + 7(x + 3)2 C. (x + 3)2 + x2 + 7x D. ln(x + 3)

OpenStudy (anonymous):

We have : \[G(F(x))=G(x^2+7x)\\ ~~~=\left(x^2+7x+3\right)^2\]